Find the z-score, to the nearest hundredth, that satisfies the given condition.

0.230 square unit of the area of the standard normal distribution is to the right of z.
Answer #1

C.230 → → P(Z > zo) = 0.230 1-P (247) = 0-230 P(2420) = 0.77 Zo = 70.77 = 0.739 90.74 t-score = 10.74 (using z table)